# Double Check My Desired Return: Transformer with Target Alignment for Offline Reinforcement Learning

### v1

@@ 4 Method
To achieve high returns, we can set an aggressive target return that even exceeds the best possible return. The model will sample actions based on this desired target and validate them with action-values, selecting the action with the highest value. Conversely, to obtain a specific moderate return, we can set that return as the target. The model will then double-check the predicted action, selecting the one with the value closest to the target return. We offer further justification of the double‐check mechanism in Appendix C.1. We include the algorithm of Doctor in Appendix B.
5 Experiments
  • What is the impact of Doctor’s double-check mechanism on target alignment, and how does it translate into improvements in the final cumulative return?
Setup
We use D4RL (Fu et al. 2020) and EpiCare (Hargrave, Spaeth, and Grosenick 2024) as our testbeds. More environmental details can be found in Appendix D.
  • D4RL. D4RL consists of various robotic environments and datasets. We consider Locomotion-v2 tasks, as well as Maze2D-v1 and Adroit-v1 tasks. To address the questions comprehensively, we test our method against 11 representative baselines, that are state of the art in each categories. For value-based methods, we select CQL (Kumar et al. 2020), which adopts pessimistic action estimation. IQL (Kostrikov, Nair, and Levine 2022), an in-sample multi-step dynamic programming approach and policy regularization-based approach, TD3+BC (Fujimoto and Gu 2021). For RvS-based methods, we consider DT (Chen et al. 2021), MTM (Wu et al. 2023), RvS-R (Emmons et al. 2021) and RADT (Tanaka et al. 2024). And we evaluate advanced methods proposed to integrate RvS and TD learning. We include CGDT (Wang et al. 2024) and ACT (Gao et al. 2024), both aiming to improve RvS-trained Transformer-based policies through value functions. QT (Hu et al. 2024) enhances the stitching ability of DT by introducing the Q-value regularizer, achieving strong performance.
  • EpiCare. EpiCare is an offline RL benchmark designed for dynamic treatment regimes (DTRs), simulating core challenges in longitudinal clinical settings. It incorporates short treatment horizons, heterogeneous treatment effects, partial observability, and adverse events. The offline dataset used for training is generated through simulated clinical trials, reflecting realistic data collection processes. In addition to cumulative returns, the benchmark also measures adverse event rates. For comparison, we adopt the methods provided by the EpiCare benchmark, BC (Torabi, Warnell, and Stone 2018), IQL (Kostrikov, Nair, and Levine 2022), TD3+BC (Fujimoto and Gu 2021), CQL (Kumar et al. 2020), and additionally implemented DT (Chen et al. 2021) and MTM (Wu et al. 2023) as our comparison baselines.
Figure 4: We evaluate the alignment ability of Doctor on the hopper-medium-replay-v2 with the top X% returns of trajectories removed. The dashed red line presents the highest return in the dataset. The dashed black lines denote the ideal alignment. Doctor achieves much better alignment across a wide range of target returns compared to DT, MTM and QT.
The Superiority of Doctor in Return Alignment
where $g_{\text{desired}}$ is the goal position, $g_{\text{achieved},t}$ is the achieved goal position at time step $t$. The horizon of an episode is 300. We evaluate Doctor under three different target returns, specifically 10, 40, and 80, depicted from left to right. The agent is guided to achieve an actual return that aligns with the target return, resulting in actual returns of 37.80, 46.85, and 80.16, respectively. We provide the complete trajectory plots of the agents in Appendix E.1.
Next, we evaluate the alignment ability of Doctor on the hopper-medium-replay-v2 dataset. We vary the level of suboptimality by removing top return trajectories from the dataset and test the model across a wide range of target returns. As suboptimality increases, the maximum returns of the trajectories in the dataset progressively decrease, moving the dataset further away from the optimal. We perform 140,000 gradient updates during training, and evaluate the model by rolling out 10 episodes. we set $N=300$ and $\delta$ is small value fluctuating based on the maximum return in the dataset. Specifically, we choose $\delta=5\%\times R_{\text{max}}$. We report the final results over five random seeds.
We compare Doctor with DT, MTM and QT, Figure 4 shows the results. The x-axis represents the target return, the y-axis represents the actual return achieved in the environment. The dashed red line marks the max return in the dataset. The dashed black lines denote the ideal line that perfectly aligns with the target return. Doctor achieves significantly better alignment, more closely adhering to the ideal diagonal line, even when the target return exceeds the maximum return in the dataset to some extent. This indicates that Doctor can not only interpolate more effectively within the dataset but also facilitates extrapolation, allowing the model to achieve target returns accurately, even beyond those observed in the training data. Notably, bidirectional transformer based methods like MTM and Doctor lie closer to the diagonal than DT and QT, highlighting the architectural advantage in terms of target alignment. We provide more results in Appendix E.2 and test the error between the actual return and the target return in Appendix E.6.
The Advantage of Doctor in Clinical Policy Control
In this section, we demonstrate how Doctor’s superior alignment can benefit medical RL applications. To this end, we evaluate Doctor using the EpiCare benchmark. EpiCare benchmark provides datasets generated by distinct behavior policies across eight different environments (defined by environment seeds 1–8, each simulating diseases with different characteristics). We use the dataset generated by the Standard of Care (SoC) policy, which simulates clinician behavior by avoiding high-risk treatments based on symptom thresholds. Baseline methods are adapted to discrete control by optimizing logits of one-hot encoded action outputs. We evaluated the models online on 1,000 trajectories, measuring key metrics: clinical returns and adverse events.
Figure 5: Mean returns and adverse‐event rates per 10k episodes under moderate (m, 0.4×) and aggressive (a, 0.8×) targets (+/-% denote change vs. BC). Results are averaged over 3 seeds. Results show that (1) DT and MTM exhibit only marginal shifts relative to BC, while Doctor demonstrates significant improvements. (2) DT and MTM have minimal responsiveness to target changes. By contrast, Doctor can sensitively adjust its policy to meet specified targets.
This result demonstrates that Doctor can flexibly control the treatment policy between aggressive and conservative as a single unified model. Such controllability holds promise for clinical practice, allowing physicians to tailor interventions to each patient’s unique condition, choosing more aggressive therapies when potential benefit outweighs risk, or adopting conservative approaches to safeguard vulnerable individuals (Rivenbark, O’Connor, and Coleman 2013). We report a comparison of Doctor against baselines on the EpiCare benchmark in terms of mean returns and adverse event rate in Appendix E.5. Doctor achieves both higher returns and lower adverse event rate, indicating its advantages in treatment efficacy and safety.
Tasks BC CQL IQL TD3+BC DT MTM RADT RvS CGDT ACT QT Doctor
[…]
Locomotion Sum 466.7 698.5 692.6 677.5 671.7 708.6 738.2 645.5 741.4 716.2 767.4 769.7
Table 1: Offline results on the D4RL benchmark. Doctor achieves competitive results compared to the baselines, including RvS-based methods and TD learning-based methods, as well as advanced methods proposed to integrate RvS and TD learning. The best baselines are highlighted in bold for final performance.
The Impact of Doctor’s Double-Check Mechanism
Figure 6: The effect of the sampling size NN. As $N$ increases, Doctor achieves better alignment with the target return.
Approach halfcheetah-M hopper-M walker2d-M
[…]
Time Complexity DT MTM QT Doctor
Inference (seconds) 0.01 0.012 0.016 0.013
Training (seconds) 2.13 1.29 2.51 1.35
Table 3: Time Complexity for Algorithms
The Capability of Doctor in Return Maximization
We further assess the return maximization performance of Doctor in offline settings, presenting results averaged over five random seeds. As shown in Table 1, Doctor achieves competitive performance compared to the strongest baseline, QT, and outperforms other baselines by a substantial margin. QT demonstrates robust performance in maximizing the return, but it degrades interpolation accuracy and fails to maintain accurate alignment, as depicted in Figure 4, whereas Doctor consistently achieves superior alignment.
We report results on the Maze2d tasks that contain very few near-optimal trajectories in Appendix E.4, Doctor outperforms other baselines, showing its stitching ability. Furthermore, in a complex high-dimensional control task, Adroit, Doctor also achieves clear improvements, as shown in Appendix E.3. These results highlight Doctor’s effectiveness in a wide range of tasks.
6 Conclusion
In this work, we propose a novel method that double checks the transformer with target alignment for offline RL (Doctor). At inference time, we introduce a double-check mechanism to sample actions based on desired target returns and validate them with action-values to enhance alignment. Experiments demonstrate that Doctor achieves superior target alignment both within and beyond the dataset. This alignment can control treatment policies that balance the adverse event rate with therapeutic returns on EpiCare benchmark.
Appendix
A Outline
In this appendix, we provide the following supporting material in order: Appendix B details the full algorithm of Doctor; Appendix C provides additional justification of double-check mechanism and further discussion about our architecture design choice; Appendix D describes the environment details; Appendix E presents extended experimental results such as visualization examples, additional return-alignment plots, additional results, absolute error estimation; and Appendix F elaborates the model and training details; Finally, we discuss the limitations and the broader impact in Appendix G and Appendix H.
B Algorithm
The algorithm of Doctor can be found in Alg. 1. Doctor repeatedly samples trajectories and updates both a self-supervised reconstruction loss and a value loss during training. In the inference phase, at each timestep, it first samples $N$ candidate returns, rolls out trajectories with the transformer, then double-checks each proposed next action by comparing its Q-value to the corresponding target return and picks the action whose value lies closest to the desired return. If online fine-tuning is enabled, actions may further be sampled from a Boltzmann distribution over Q-values to encourage exploration.
@@ B Algorithm
  if online fine-tuning then
   Select action from Boltzmann distribution via $\pi\left(a_{t}\mid s_{t}\right)=\frac{\exp\left(\beta q_{t,i}\right)}{\sum_{i}\exp\left(\beta q_{t,i}\right)}$.
  end if
@@
C Discussion
C.1 Additional Justification of Double Check Mechanism
Assume our action-value function is optimal, $Q=Q^{*}$. We select the action $a_{t}$ that minimizes the absolute difference between the predicted action-value and the desired return:
$a_{t}=\arg\min_{a}\left|Q^{*}\left(s_{t},a\right)-R_{t}\right|.$ (11)
We show that this action selection aligns the expected return with $R_{t}$, achieving return alignment.
@@ C.1 Additional Justification of Double Check Mechanism
C.2 Additional Explanation for the Design Choice
@@
Transformer Architecture. Early works like DT (Chen et al. 2021) and TT (Janner, Li, and Levine 2021) rely on causal transformers, which predict each subsequent token based on past tokens. Recent studies such as MTM (Wu et al. 2023) demonstrated that bidirectional transformers trained with a mixture of random and auto-regressive masking achieve good performance. Random masks encourage the model to reconstruct missing tokens by leveraging the context of the unmasked sequence, enhancing the reconstruction and generalization abilities. Our experiments also demonstrate that the bidirectional-based MTM outperforms DT, exhibiting stronger interpolation when handling trajectories associated with underrepresented returns. Our proposed model further enhances these benefits, resulting in improved return alignment.
@@
Q-function Learning. To learn the Q-function, we employ expectile regression (Kostrikov, Nair, and Levine 2022). When the expectile parameter $\nu>0.5$, the resulting loss function assigns lower weight to errors arising from underestimated Q-values. This asymmetric allows us to approximate $Q(s,a)\leftarrow r+\gamma\max_{\left(s^{\prime},a^{\prime}\right)\in\mathcal{D}}Q\left(s^{\prime},a^{\prime}\right)$, achieving reliable in-sample learning. This precision is especially critical for our double-check mechanism, which is dependent on accurate Q-value predictions within data-supported regions, selecting actions closely aligned with the target return.
C.3 Difference Between Doctor and Previous Methods
RvS fails to achieve one of the desired properties of offline RL agents, stitching. To enhance transformers with stitching ability, QDT (Yamagata, Khalil, and Santos-Rodriguez 2023) relabels the Return-to-Go in the dataset by learning a conservative value function, serves as data augmentation by incorporating stitched trajectories into the training dataset. EDT (Wu, Wang, and Hamaya 2023) optimizes the trajectory by retaining a longer history when the previous trajectory is optimal and a shorter one when it is sub-optimal, enabling it to stitch with a more optimal trajectory. ACT (Gao et al. 2024) and CGDT (Wang et al. 2024) uses learned Q-values to guide actions or as critic prompts. QT (Hu et al. 2024) and QCS (Kim et al. 2024) impose a regularizer of the Q-value during transformer training.
Different from these previous methods, which mainly focus on improving the model’s stitching ability to obtain the best policy from the dataset. Our primary objective is explicit target alignment, enabling a single offline trained transformer to perform accurately at different policy expertise. Technically, unlike previous methods that learn the value functions beforehand and train the SL model separately afterward, Doctor let the value function utilize the representation feature of the transformer base model. And we perform test-time scaling of the outputs, combined with value function verification to ensure alignment.
D Environment Details
D.1 D4RL
@@
Gym Locomotion. The D4RL Gym locomotion benchmark  (Fu et al. 2020) includes environments provided by OpenAI Gym (Brockman 2016), specifically Walker2d, Hopper, and HalfCheetah. Each task is configured with three levels of dataset difficulty: Medium-Replay, Medium, and Medium-Expert. The Medium dataset corresponds to policies performing at approximately one-third of expert-level performance, while the Medium-Replay dataset contains the replay buffer from an agent trained to medium-level performance. The Medium-Expert dataset combines trajectories generated by both medium and expert policies. These environments are widely used for evaluating reinforcement learning algorithms. For expample, Walker2d environment simulates a robot tasked with walking as fast and as stably as possible. The robot must coordinate its two legs to achieve efficient locomotion without falling over. These environments are designed to test an agent’s ability to learn complex motor skills and optimize control strategies.
Adroit (Rajeswaran et al. 2017). Adroit is a suite of dexterous manipulation tasks designed to simulate the control of a five-fingered robotic hand. Our experiments focus on three tasks from this suite: Pen, Door, and Hammer. For example, the Pen task involves rotating a pen to a specific orientation using the robotic hand’s dexterous manipulation skills. The cloned tasks used in our experiment collect a 50-50 split of demonstration data and 2500 trajectories sampled from a behavior cloning policy.
@@
Maze2D. Maze2D is a navigation task in which the agent is required to reach a fixed target position. These tasks are designed to evaluate the ability of offline reinforcement learning algorithms to stitch together different trajectory fragments  (Fu et al. 2020). We used three environments included in Maze2D, umaze, medium, and large, with complexity and path length to the target increasing sequentially.
D.2 EpiCare

### v2

@@ 4.3 Inference Time Alignment
To achieve high returns, we can set an aggressive target return that even exceeds the best possible return. The model will sample actions based on this desired target and validate them with action-values, selecting the action with the highest value. Conversely, to obtain a specific moderate return, we can set that return as the target. The model will then double-check the predicted action, selecting the one with the value closest to the target return. We include the algorithm of Doctor in Alg. 1.
@@
4.4 Justification of Double Check Mechanism
To better understand why the proposed double-check mechanism leads to accurate target alignment, we provide a further justification in this section.
Assume our action-value function is optimal, $Q=Q^{*}$. We select the action $a_{t}$ that minimizes the absolute difference between the predicted action-value and the desired return according to $a_{t}=\arg\min_{a}\left|Q^{*}\left(s_{t},a\right)-R_{t}\right|.$ We show that this action selection aligns the expected return with $R_{t}$, achieving return alignment.
@@ 4.4 Justification of Double Check Mechanism
5 Experiments
  • 1.
  • 2.
    What is the impact of Doctor’s double-check mechanism on target alignment, and how does it translate into improvements in the final performance?
  • 3.
5.1 Setup
  • 1.
    Benchmarks. We use D4RL [15] and EpiCare [5] as our testbeds. D4RL consists of various robotic control environments and datasets. We consider Locomotion-v2 tasks, as well as Maze2D-v1 and Adroit-v1 tasks on the D4RL benchmark. EpiCare is an offline RL benchmark designed for dynamic treatment regimes (DTRs), simulating core challenges in longitudinal clinical settings. It incorporates short treatment horizons, heterogeneous treatment effects, partial observability, and adverse events. The offline dataset used for training is generated through simulated clinical trials, reflecting realistic data collection processes. More environmental details can be found in C.
  • 2.
    Baselines. To address the questions comprehensively, we test our method against representative baselines that are state of the art in each category. We select CQL [23], which adopts pessimistic action estimation. IQL [18], an in-sample multi-step dynamic programming approach. We also include policy regularization-based approaches like TD3+BC [20] as comparison baselines. For RvS-based approaches, we include BC, which maximizes the likelihood of the dataset. DT [12], a transformer-based return conditioned BC, which is shown to be effective on D4RL. And MTM [27] employs a combination of random and autoregressive mask patterns.
Figure 4: We evaluate the alignment ability of Doctor on hopper-medium-replay-v2 (top) and walker2d-medium-replay-v2 (bottom) tasks with the top X% returns of trajectories removed. The dashed red line presents the highest return in the dataset. The dashed black lines denote the ideal alignment. Doctor achieves much better alignment across a wide range of target returns compared to DT, MTM.
5.2 The Superiority of Doctor in Return Alignment
where $g_{\text{desired}}$ is the goal position, $g_{\text{achieved},t}$ is the achieved goal position at time step $t$. The horizon of an episode is 300. We evaluate Doctor under three different target returns, specifically 10, 40, and 80, depicted from left to right. The agent is guided to achieve an actual return that aligns with the target return, resulting in actual returns of 37.80, 46.85, and 80.16, respectively. We provide the complete trajectory plots of the agents in D.
Next, we evaluate the alignment ability of Doctor on the hopper-medium-replay-v2 and walker2d-medium-replay-v2 datasets. We vary the level of suboptimality by removing top return trajectories from the dataset and test the model across a wide range of target returns. As suboptimality increases, the maximum returns of the trajectories in the dataset progressively decrease, moving the dataset further away from the optimal.
Figure 4 shows the results. The x-axis represents the target return, the y-axis represents the actual return achieved in the environment. The dashed red line marks the max return in the dataset. The dashed black lines denote the ideal line that perfectly aligns with the target return. Doctor achieves significantly better alignment, more closely adhering to the ideal diagonal line, even when the target return exceeds the maximum return in the dataset to some extent. This indicates that Doctor can not only interpolate more effectively within the dataset but also facilitates extrapolation, allowing the model to achieve target returns accurately, even beyond those observed in the training data. Notably, bidirectional transformer-based methods like MTM and Doctor lie closer to the diagonal than DT, highlighting the architectural advantage in terms of target alignment.
5.3 The Impact of Doctor’s Double-Check Mechanism
Figure 5: The effect of the sample number NN in Doctor. As $N$ increases, Doctor achieves better alignment with the target return.
Time Complexity DT MTM Doctor
Inference (seconds) 0.01 0.012 0.013
Training (seconds) 2.13 1.29 1.35
Table 1: Time Complexity for Algorithms
Approach halfcheetah-M hopper-M walker2d-M
[…]
Locomotion BC CQL IQL TD3+BC DT MTM Doctor
[…]
Sum 472.6 694.8 717.8 696.9 676.8 714.3 769.7
Table 3: Offline results on the D4RL benchmark. Doctor achieves competitive results compared to the baselines, including RvS-based methods and TD learning-based methods. The best baselines are highlighted in bold for final performance.
5.4 The Capability of Doctor in Return Maximization
We further assess the return maximization performance of Doctor in offline settings, presenting results averaged over five random seeds. As shown in Table 3, Doctor outperforms other baselines by a substantial margin and achieves the strongest results in 8 out of 9 tasks on Gym locomotion tasks. Doctor integrates supervised learning with TD learning, allowing the model to benefit from the strengths of both paradigms. In datasets with expert-level trajectories such as medium-expert (ME) tasks in the Locomotion environment, Doctor performs better than TD learning-based methods like IQL and CQL, demonstrating its ability of sequence modeling. In datasets with medium-level trajectories such as Medium and Medium-Replay, Doctor outperforms RvS-based methods like MTM, and DT, which indicates the stitching capability due to the value functions. This suggests that Doctor effectively integrates the advantages of both approaches, achieving superior performance with both low-return and high-return datasets.
Adroit BC CQL IQL TD3+BC DT MTM Doctor
[…]
Sum 56.6 7.8 88.3 5.4 75.1 96.5 100.2
Table 4: Offline results on the Adroit-cloned tasks. Doctor achieves competitive or superior results compared to the baselines.
Maze2D BC CQL IQL TD3+BC DT MTM Doctor
[…]
Sum 18.4 228.5 140.0 189.5 77.9 89.0 279.1
Table 5: Offline results on the maze2d tasks. Doctor achieves competitive or superior results compared to the baselines.
In addition to evaluation on the locomotion tasks, we further assess Doctor in the challenging Adroit tasks, which consist of three dexterous control tasks, each requiring precise interactions and hand-object coordination. As shown in Table 4, Doctor outperforms other baselines on the Pen environment, demonstrating Doctor’s robustness in these substantially harder high-dimensional dexterous manipulation tasks.
These results highlight Doctor’s effectiveness in a wide range of tasks.
5.5 The Advantage of Doctor in Clinical Policy Control
In this section, we demonstrate how Doctor’s superior alignment can benefit RL applications. To this end, we evaluate Doctor using the EpiCare benchmark. EpiCare benchmark provides datasets generated by distinct behavior policies across eight different environments (defined by environment seeds 1–8, each simulating diseases with different characteristics). We use the dataset generated by the Standard of Care (SoC) policy, which simulates clinician behavior by avoiding high-risk treatments based on symptom thresholds. Baseline methods are adapted to discrete control by optimizing logits of one-hot encoded action outputs. We evaluated the models online on 1,000 trajectories, measuring key metrics: clinical returns and adverse events. Clinical returns quantify the overall therapeutic benefit of the treatment policy, while adverse events capture the frequency of unsafe outcomes.
Figure 6: Mean returns and adverse‐event rates per 10k episodes under moderate (m, 0.4×) and aggressive (a, 0.8×) targets. Numbers denote improvement or decrease compared with BC. Results show that (1) DT and MTM exhibit only marginal shifts relative to BC, while Doctor demonstrates significant improvements. (2) DT and MTM have minimal responsiveness to target changes. By contrast, Doctor can sensitively adjust its policy to meet specified targets.
This result demonstrates that Doctor can flexibly control the treatment policy between aggressive and conservative as a single unified model. Such controllability holds promise for clinical practice, allowing physicians to tailor interventions to each patient’s unique condition, choosing more aggressive therapies when potential benefit outweighs risk, or adopting conservative approaches to safeguard vulnerable individuals.
We further report a comparison of Doctor against baselines on the EpiCare benchmark in terms of mean returns and adverse event rate on the SoC dataset, as shown in Figure 7. Doctor achieves the lowest adverse events and the best clinical returns, outperforming other baselines and demonstrating that Doctor’s alignment produces substantially safer and more effective treatment policies on the dynamic treatment regime benchmark.
Figure 7: Mean returns and number of adverse events per 10k episodes (lower is better). Results are averaged over eight structurally distinct EpiCare environments (environment seeds 1–8) for each offline learning method on the SoC dataset. Error bars denote standard deviation.
6 Conclusion
This work introduced Doctor, a novel offline reinforcement learning approach that integrates a double-check mechanism to enhance target alignment in return-conditioned transformers. By jointly optimizing supervised trajectory reconstruction and temporal-difference value estimation, Doctor demonstrates the ability to interpolate across underrepresented return ranges and extrapolate beyond dataset-supported regions. Experimental results on diverse benchmarks, including D4RL and EpiCare, confirm its superiority over existing methods in both return maximization and controllable policy modulation. These findings highlight the strengths of Doctor in balancing therapeutic efficacy with safety in clinical decision-making and in offering fine-grained control in robotics.
At the same time, it is important to recognize the limitations of our study. The evaluation of Doctor on medical decision-making tasks remains in a simulated environment, which, although captures many real-world clinical challenges, such as short treatment horizons, heterogeneous treatment effects, partial observability and adverse events, may not fully reflect real-world healthcare data. As a next step, we plan to evaluate our method on real clinical datasets.
Beyond these strengths and weaknesses, our contributions may also benefit current fields. This study focuses on improving the alignment accuracy of RvS methods in offline reinforcement learning. By addressing their current limitations and achieving a closer match between actual and target returns, it advances the field of offline reinforcement learning. The general offline reinforcement learning community may benefit from adopting Doctor as a robust baseline. In robotic control, precise alignment between policy outputs and performance targets could enable safer and more reliable manipulation. In healthcare, adaptive treatment planning based on target-aligned policies could support personalized medicine by balancing efficacy and risk. As foundational machine learning research, it poses no negative societal impacts.
Finally, we see promising avenues for future work. A particularly exciting direction is to explore the integration of the double-check mechanism with large foundation models. Recent advances in generative transformers and multimodal learning provide opportunities to incorporate Doctor’s target alignment into broader architectures, enabling alignment of outputs with diverse user-defined targets across tasks and modalities. This could, for example, support the development of controllable base models that meet safety, fairness, and performance objectives. Investigating such integration, alongside applying the method to real-world applications, constitutes a vital step for future research.
Appendix A Outline
In this appendix, we provide the following supporting material in order: Appendix B provides further discussion about our architecture design choice; Appendix C describes the environment details; Appendix D presents extended experimental results such as visualization examples and additional results; and Appendix E offers the model and training details.
Appendix B Additional Discussion for the Design Choice
@@
Transformer Architecture. Early works like DT [12] and TT [13] rely on causal transformers, which predict each subsequent token based on past tokens. Recent studies such as MTM [27] demonstrated that bidirectional transformers trained with a mixture of random and auto-regressive masking achieve good performance. Random masks encourage the model to reconstruct missing tokens by leveraging the context of the unmasked sequence, enhancing the reconstruction and generalization abilities. Our experiments also demonstrate that the bidirectional-based MTM outperforms DT, exhibiting stronger interpolation when handling trajectories associated with underrepresented returns. Our proposed model further enhances these benefits, resulting in improved return alignment.
@@
Q-function Learning. To learn the Q-function, we employ expectile regression [18]. When the expectile parameter $\nu>0.5$, the resulting loss function assigns lower weight to errors arising from underestimated Q-values. This asymmetric allows us to approximate $Q(s,a)\leftarrow r+\gamma\max_{\left(s^{\prime},a^{\prime}\right)\in\mathcal{D}}Q\left(s^{\prime},a^{\prime}\right)$, achieving reliable in-sample learning. This precision is especially critical for our double-check mechanism, which is dependent on accurate Q-value predictions within data-supported regions, selecting actions closely aligned with the target return.
Appendix C Environment Details
C.1 D4RL
@@
Gym Locomotion. The D4RL Gym locomotion benchmark  [15] includes environments provided by OpenAI Gym, specifically Walker2d, Hopper, and HalfCheetah. Each task is configured with three levels of dataset difficulty: Medium-Replay, Medium, and Medium-Expert. The Medium dataset corresponds to policies performing at approximately one-third of expert-level performance, while the Medium-Replay dataset contains the replay buffer from an agent trained to medium-level performance. The Medium-Expert dataset combines trajectories generated by both medium and expert policies. These environments are widely used for evaluating reinforcement learning algorithms. For expample, Walker2d environment simulates a robot tasked with walking as fast and as stably as possible. The robot must coordinate its two legs to achieve efficient locomotion without falling over. These environments are designed to test an agent’s ability to learn complex motor skills and optimize control strategies. Adroit. Adroit is a suite of dexterous manipulation tasks designed to simulate the control of a five-fingered robotic hand. Our experiments focus on three tasks from this suite: Pen, Door, and Hammer. For example, the Pen task involves rotating a pen to a specific orientation using the robotic hand’s dexterous manipulation skills. The cloned tasks used in our experiment collect a 50-50 split of demonstration data and 2500 trajectories sampled from a behavior cloning policy.
@@
Maze2D. Maze2D is a navigation task in which the agent is required to reach a fixed target position. These tasks are designed to evaluate the ability of offline reinforcement learning algorithms to stitch together different trajectory fragments  [15]. We used three environments included in Maze2D, umaze, medium, and large, with complexity and path length to the target increasing sequentially.
C.2 EpiCare